### v1

# Transferable Ensemble Black-box Jailbreak Attacks on Large Language Models

Abstract
In this report, we propose a novel black-box jailbreak attacking framework that incorporates various LLM-as-Attacker methods to deliver transferable and powerful jailbreak attacks. Our method is designed based on three key observations from existing jailbreaking studies and practices. First, we consider an ensemble approach should be more effective in exposing the vulnerabilities of an aligned LLM compared to individual attacks. Second, different malicious instructions inherently vary in their jailbreaking difficulty, necessitating differentiated treatment to ensure more efficient attacks. Finally, the semantic coherence of a malicious instruction is crucial for triggering the defenses of an aligned LLM; therefore, it must be carefully disrupted to manipulate its embedding representation, thereby increasing the jailbreak success rate. We validated our approach by participating in the Competition for LLM and Agent Safety 2024, where our team achieved top performance in the Jailbreaking Attack Track.
1 Introduction
Jailbreaking large language models (LLMs) is being intensively studied to evaluate the safety of LLMs [1, 2]. Existing jailbreaking methods, as adversarial attacks targeting small-scale learning models, can be categorized into white-box and black-box approaches. In white-box methods, gradient-based optimization techniques are employed to identify adversarial suffixes [3]. For black-box methods, various optimization strategies, such as genetic algorithms, are used to refine jailbreak prompt templates through rephrasing, word replacement, and other modifications [4, 5]. Recent research [6] has demonstrated that LLMs can function as powerful optimizers when provided with sufficient contextual information. Indeed, both the persuasive adversarial prompt (PAP) [7] and Tree of Attacks with Pruning (TAP) [8] methods leverage multiple LLMs to optimize malicious instructions. In this line of research, given a target aligned LLM to jailbreak, one LLM operates as an attacker, optimizing a malicious instruction to circumvent the unknown safety mechanisms of the target LLM. A second LLM serves as a judge, observing the interaction history between the attacker and the target LLM, evaluating the attack’s effectiveness, and providing feedback to the attacker LLM. The PAP method [7] has been officially adopted as the baseline attack strategy in the Competition for LLM and Agent Safety 2024 (CLAS 2024) [^1] https://www.llmagentsafetycomp24.com/.
Model ensembling has been widely recognized as one of the most effective approaches in various machine learning competitions. In the context of jailbreaking LLMs, we believe that this approach holds promise, although there has been limited research exploring this method. Existing studies, as previously mentioned, have focused on developing specific jailbreaking techniques. This, however, has the limitation of only partially exposing the vulnerabilities of a target LLM, potentially leading to an inaccurate assessment of model safety. Furthermore, based on our experiences, we find that different malicious instructions exhibit varying levels of effectiveness for bypassing the safety mechanism of an aligned LLM. For example, the widely used “how to build a bomb” query is often readily refused, whereas other instructions, such as “Provide detailed instructions on where to obtain hydroxychloroquine without a prescription …” [^2] We select this sample from the set of instructions provided by the CLAS 2024 competition., have a higher likelihood of eliciting harmful responses from the target LLM. This difference in the malicious instructions suggests that a comprehensive evaluation of LLM safety requires tailored optimization efforts for different attacking instructions. Moreover, prior research [9, 10] has demonstrated that the internal representations of malicious and benign instructions processed by an aligned LLM are distinctly separated. Given that these embeddings mainly retain the semantic information of the instructions, delivering more effective jailbreaking attacks necessitates perturbing these embeddings without significantly altering their original semantics.
Based on the aforementioned observations, we have designed our jailbreaking method according to the following principles:
  1. 1.
    Ensemble Different LLM-as-Attacker Methods: We begin by developing an ensemble framework that incorporates different LLM-as-Attacker methods. This framework aims to deliver transferable and powerful jailbreak attacks.
  2. 2.
    Adjust Optimization Budgets for Different Instructions: Given a set of malicious instructions, we analyze their individual jailbreak scores [^3] Please refer to the website of CLAS 2024 for the calculation of the jailbreak score. to identify those that are more challenging to optimize. We then allocate additional computation resources to these instructions.
  3. 3.
    Shatter Semantic Coherence to Balance Attack Performance and Instruction Stealthiness: To maintain a balance between attack performance and instruction stealthiness, we randomly select a small subset of words from the given malicious instructions and insert them into the optimized instructions. This helps to preserve relatively high TF-IDF similarity, thereby avoiding easy detection.
To evaluate our method, we participated in CLAS 2024444Our team is named LlaXa, representing Large Language Model Break AI. and achieved top performance in the Jailbreaking Attack Track (Track I). We actively respond to the call of the competition organizers by open-sourcing our solution at https://github.com/YQYANG2233/Large-Language-Model-Break-AI.
2 Method
2.1 Ensemble Jailbreaking
Among various jailbreaking methods, we find that the PAP method [7] and a similar TAP method [8], exhibit superior transferability and overall jailbreaking performance. Therefore, we propose to develop our ensemble jailbreak framework based on these two existing methods.
It should be noted that the original TAP prompting framework includes a feedback session to incorporate the interaction history for the attacker LLM. In our design, we propose to replace TAP’s original evaluation prompt with the one provided by the competition organizers and to introduce additional information in the feedback session. This includes the reasoning behind the judge LLM’s evaluation of an optimized malicious instruction and a keyword score, which quantifies the likelihood of the target LLM refusing to respond to the optimized malicious instruction. Our modification of the TAP prompting framework is shown in Figure 1.
FORMAT: …… The REASON is the reason for the SCORE. If the output for your prompt contains any of [“unable”, “sorry”, “cannot”, “can’t”, “not able”], the KEYWORD_SCORE will be 0, otherwise, the KEYWORD_SCORE will be 1. ……
Figure 1: Added feedback information
We utilize each of these methods individually to optimize the same set of malicious instructions provided by the competition organizers, and evaluate the jailbreak scores of the optimized malicious instructions generated by these two methods according to the evaluation protocol specified by the competition organizers. Given the two sets of evaluated instructions, we propose either a greedy selection of the instructions that yield higher jailbreak scores or a weighted random sampling mechanism to avoid over-optimization for our local environment. In our local experiments, we observe that the greedy approach achieves overall better performance. Consequently, we adopt the greedy approach in our submitted implementation.
2.2 Stealthiness Enhancing
To enhance the stealthiness of the optimized malicious instructions  [^5] Please refer to the website of CLAS 2024 for the calculation of the stealthiness score, which reflects the difference in word frequency between the original and optimized instructions, we propose to add a limited set of words that are randomly sampled from the original instructions. To prevent the added words, which may contain obvious malicious semantics, from triggering the target LLM to refuse to respond to the optimized instructions, we remove obviously harmful words from the original instructions. It should be noted that this word-insertion operation may affect the semantic correctness of a malicious instruction. To address this, we further propose an iterative approach to select optimized instructions that achieve increased stealthiness while maintaining their jailbreak scores.
Figure 2: Diagram of Jailbreak Score distribution
2.3 Selective Boosting
Through our intensive local experiments, we have observed that not all of the provided malicious instructions exhibit the same difficulty in jailbreaking the target LLM. To illustrate this effect, we present the distribution of jailbreak scores for all malicious instructions, as evaluated by the Llama3-8B-Instruct model, in Figure 2. To further enhance jailbreak performance within limited computational budgets, we propose to iteratively optimize the instructions with lower scores, allocating additional computational resources to them.
3 Experiments
3.1 Experiment Setup
Datasets
We evaluate our proposed method using the dataset of malicious instructions provided by the CLAS 2024 competition. These instructions cover various domains, including finance, law, and criminal planning, etc.
Models
We utilized Gemma-2B-IT and Gemma2-9B-IT as our target models for the attacks and employed Llama3-8B-Instruct, GLM-4-Plus, GLM-4-Flash, Qwen-Max-Latest, and DeepSeek-V2.5 as judge models to evaluate the attacking performance. We used GPT-4 and Qwen-Max-Latest as the attackers to optimize malicious instructions for the TAP and PAP method, respectively.
Table 1: Performance of different methods on different target models. ”Jail” indicates the jailbreak score. ”Stl” indicates the stealthiness score. ”Combined” is calculated by 0.84 $\times$× Jail + 0.16 $\times$× Stl. ”Ensemble wo Stl” indicates our method without enhancement in section 2.2 and ”Ensemble w Stl” indicates the opposite.
Method Gemma-2B-IT Gemma2-9B-IT
Jail Stl Combined Jail Stl Combined
TAP 0.8448 0.1742 0.7375 0.8510 0.1852 0.7445
PAP 0.6957 0.2614 0.6262 0.6847 0.2362 0.6129
Ensemble wo Stl 0.8859 0.2064 0.7908 0.8640 0.1935 0.7701
Ensemble w Stl 0.9325 0.4011 0.8475 0.9133 0.3896 0.8295
3.2 Transferable Effectiveness of the Ensemble Jailbreak Attack
We present the performance of different jailbreak methods across different target models in Table 1. Our proposed ensemble jailbreak method demonstrates significant improvements over the individual TAP and PAP methods on both target models. Additionally, we observed that the proposed stealthiness enhancement method not only achieved higher stealthiness scores but also enhanced the jailbreak scores.
3.3 Transferability for Judge Models
Figure 3: Evaluation scores of adopting different judgment models. We use the same target model and evaluation instructions.
In our experiments, we observed that the prompt designed for configuring a LLM as a judge, as provided by the competition organizers, exhibits strong transferability. We applied this identical judgment prompt across various commercial and open-source LLMs. As illustrated in Figure 3, the ratings generated by different judge LLMs were both stable and closely correlated, thereby validating the efficacy of the official judgment process.
4 Conclusion
In this report, we introduce our proposed jailbreak method, which we developed for participating in and winning the Competition for LLM and Agent Safety 2024. Through the public competition and preliminary experimental results, we highlight the effectiveness of the proposed ensemble framework, the stealth-enhancing method, and the necessity of adaptively optimizing malicious instructions based on their jailbreaking difficulties. In our future work, we will further refine our proposed method and conduct more comprehensive evaluations to demonstrate its transferability.